OpenStudy (anonymous):

7x/11 divided by 14xcubed/66y

OpenStudy (jdoe0001):

\(\bf \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\\ \quad \\ \quad \\ \cfrac{7x}{11}\div \cfrac{14x^3}{66y}\implies \cfrac{\frac{7x}{11}}{\frac{14x^3}{66y}}\implies \cfrac{7x}{11}\cdot \cfrac{66y}{14x^3}\)

OpenStudy (anonymous):

how do I get it to lowest terms @jdoe0001

OpenStudy (jdoe0001):

\(\bf \cfrac{7x}{11}\cdot \cfrac{66y}{14x^3}\implies \cfrac{7x}{11}\cdot \cfrac{11\cdot 6y}{7\cdot 2\cdot x\cdot x^2}\) can you cancel anything out?

OpenStudy (jdoe0001):

well... did you cancel anything out? does it have anything we can cancel out? what do you think?

OpenStudy (anonymous):

the x

OpenStudy (anonymous):

all the like terms are going to cancel out so I am left with 6y/2x squared

OpenStudy (jdoe0001):

lemme rewrite that some ...

OpenStudy (jdoe0001):

\(\bf \cfrac{7x}{11}\cdot \cfrac{66y}{14x^3}\implies \cfrac{7x}{11}\cdot \cfrac{11\cdot 6y}{7\cdot 2\cdot x\cdot x^2}\implies \cfrac{{\color{blue}{ 7}}\cdot {\color{blue}{ x}}}{{\color{blue}{ 11}}}\cdot \cfrac{{\color{blue}{ 11}}\cdot 3\cdot {\color{blue}{ 2}}\cdot y}{{\color{blue}{ 7}}\cdot {\color{blue}{ 2}}\cdot {\color{blue}{ x}}\cdot x^2}\) what do you think now, anything we can cancel out?

OpenStudy (anonymous):

After I cancel out the like terms I am left with 3y/x squared

OpenStudy (jdoe0001):

\(\bf \cfrac{\cancel{7x}}{\cancel{11}}\cdot \cfrac{\cancel{11}\cdot 3\cdot \cancel{2}y}{\cancel{7}\cdot \cancel{2}\cdot \cancel{x}\cdot x^2}\) After I cancel out the like terms I am left with 3y/x squared \(\large \checkmark\)

OpenStudy (anonymous):

\[3y \square-12y/9y cubed\]